FIND THE VALUE OF [1/2]+[1/2+1/1000]+...........[1/2+2946/1000]


[.] denotes for GIF......GREATEST INTEGER FACTOR ?

Grade:12

1 Answers

AKASH GOYAL AskiitiansExpert-IITD
13 years ago

Dear Manoj

We have to divide the series in parts

  1. [1/2]+[1/2 + 1/1000] +.................+ [1/2 + 499/1000]=0
  2. [1/2 + 500/1000]+...........+ [1/2 + 1499/1000] = 1000 * 1=1000
  3. [1/2 + 1500/1000]+...........+ [1/2 + 2499/1000] = 1000 * 2=2000
  4. [1/2 + 2500/1000]+...........+ [1/2 + 2946/1000] = 447 * 3=1341

adding all

0 + 1000 + 2000 + 1341 = 4341 Answer

Please feel free to post as many doubts on our discussion forum as you can. We are all IITians and here to help you in your IIT JEE preparation.


Now you can win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.


AKASH GOYAL
AskiitiansExpert-IITD